write an equation of line ab in slope-intercept form
a- (-4,-3)
b- (2,1)

Answer by mananth(16946)

x1 y1 x2 y2
-4 -3 2 1

slope m = (y2-y1)/(x2-x1)
( 1 - -3 )/( 2 - -4 )
( 4 / 6 )
m= 2/ 3

Plug value of the slope and point ( 2 , 1 ) in
Y = m x + b
1.00 = 1 1/3 + b
b= 1 - 1 1/3
b= - 1/ 3
So the equation will be
Y = 2/3 x - 1/3
m.ananth@hotmail.ca